OpenStudy (anonymous):

How many moles of sodium are in 100 mL of a 0.50 M solution of Na2SO4?

OpenStudy (anonymous):

mole = volume x concentration mole = (100/1000)dm^3 x 0.5M = 0.05 mol

sam (.sam.):

This isn't complete yet, 0.05moles are in \(Na_2SO_4\). You have to find out the moles of sodium, Na, in \(Na_2SO_4\), that is: \[0.05molNa_2SO_4 \times \frac{2~molNa}{1~molNa_2SO_4}=0.1~mol~of~Na\]